# A Novel Hybrid Approach to Contraceptive Demand Forecasting: Integrating Point Predictions with Probabilistic Distributions

### v1

@@ 1 Introduction
To our knowledge, no previous work has focused on probabilistic forecasting in contraceptive demand estimation within the FPSC at the local healthcare site level. Thus, this paper first addresses this gap by investigating a probabilistic forecasting approach for estimating demand for contraceptive products using data from January 2016 to December 2019, extracted from the Logistics Management Information System (LMIS) of Cote d’Ivoire. This is the same dataset that was used in the competition. Moreover, since the publication of the Contraceptive Forecasting Handbook, which focuses on simple forecasting methods (see USAID [62] for more information), there has been no literature evaluating the applicability and usability of different forecasting methods in contraceptive demand forecasting. Therefore, in this study, we employ a range of forecasting methods, including time series, Bayesian, ML and foundational time series methods, to produce point forecasts along with probabilistic forecasts for all products across all healthcare sites.
@@
Given that system-generated forecasts and human forecasts offer distinct benefits, it is vital to design a hybrid intelligence system that combines them. In this context, where site-level demand planners produce point forecasts, we are particularly interested in how to combine point forecasts with probabilistic forecasts to produce combined probabilistic forecasts. However, the literature often treats point forecast combination methods and probabilistic forecast combination methods separately [66]. To address this gap, we propose a Constrained Quantile Regression Averaging (CQRA) method to combine point forecasts made by experts with probabilistic forecasts generated by a system-based forecasting method. We compare the forecast performance using the Mean Absolute Scaled Error (MASE), a scale-independent metric that provides robustness and stability for point forecasts, and Continuous Ranked Probability Scores (CRPS), a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution for probabilistic forecasts in a cross-validation setup. Finally, we compare our method results against the submissions from the competition.
@@ 2.2 Human judgment in contraceptive demand forecasting
Several methods exist for combining human judgment with algorithmic forecasts. One approach involves incorporating human expertise as a feature or variable within the forecasting algorithm itself [18]. Alternatively, experts can adjust the forecasts generated by algorithms to reflect their contextual knowledge and insights [21]. Another method is to produce a purely judgmental forecast separately and then combine it with the algorithm-based forecasts [25]. Each of these strategies aims to leverage the strengths of both human expertise and algorithmic precision to enhance the overall accuracy of forecasting methods.
@@ 2.3 Literature limitations summary
These gaps in the literature and competition, along with the need for improved contraceptive demand forecasting, motivate our research to explore and assess various forecasting approaches. The key limitations identified are summarized below:
  • 1.
    Current approaches largely ignore probabilistic forecasting in contraceptive demand forecasting, leaving a gap in representing uncertainty for future demand. There is a need to develop and apply probabilistic forecasting models to better account for and manage uncertainty.
  • 2.
    Demand variations at the local level are unique and not adequately addressed by current methods, which often rely on simplistic estimation approaches. There is a need for reliable forecasting methods that can handle these variations more effectively. There is also a lack of empirical assessment of different forecasting methods for contraceptive demand forecasting.
  • 3.
    Another limitation is that there is a limited literature on how to combine point forecasts with probabilistic forecasts. This presents an opportunity to explore and develop methods that combine these approaches to enhance forecasting accuracy and reliability in the FPSC.
@@ 3 Proposed hybrid approach
We propose a CQRA model to generate a combined probabilistic forecast, utilizing both point and probabilistic forecasts. This approach builds upon the CQRA model introduced by Wang et al. [67], which focuses on combining multiple probabilistic forecasts to produce a consolidated forecast distribution. The key concept in our proposed method is to generate quantiles from a given probabilistic forecast and adjust each quantile using weights. These weights are determined by treating the point forecast as the “new reality” and formulating a linear programming (LP) problem that minimizes both the pinball loss and the absolute error between point forecast and mean of the weighted quantile forecast. The pinball loss is a strictly proper scoring rule used to evaluate quantile forecasts. It measures overall quantile performance by rewarding sharpness and penalizing miscalibration [29].
@@
However, probabilistic forecasts may have a more dominant influence in this approach since the weights $w_{i}$ are restricted to the range $[0,1]$. This implies that point forecasts are expected to align closely with the forecast distribution. When the probabilistic forecasts are reliable and the point predictions do not significantly deviate from the mean of the probabilistic forecast, this approach is recommended. However, recognizing that this is often not the case in practice, we modify the weights by allowing them to exceed 1, and we introduce a bias term to the final adjusted quantile forecast.
@@ 4.2 Data exploration
Figure 4: The trend strength and seasonality in the time series of stock distributed. The scatter plot comprises 1360 data points, with each point representing a specific time series.
@@ 4.3 Forecasting setup
In our study, we applied cross-validation, using one-step-ahead rolling windows with a 3-step-ahead forecasting horizon to align with the competition requirements. We limited the number of rolling origins to 3 per series due to computational constraints, but this still provided us with meaningful insights into method performance over time. For forecasting, we employed recursive multi-step forecasting and we generated 1000 future paths per a series. All method development and hyper-parameter tuning were conducted using only the training data to ensure that the evaluation remained unbiased and the methods were properly validated.
@@ 4.5 Performance evaluation
To assess the performance of our forecasting methods, we used both point forecast and probabilistic forecast evaluation metrics. We evaluated point forecasts using the MASE, a scale-independent metric that provides robustness, and stability [35]. The MASE formula is:
@@
To evaluate the accuracy of probabilistic forecasts, we employed the CRPS, a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution. The CRPS is given by:
@@ 5 Analysis and results
Figure 5: Average ranks and 95% confidence intervals for all the forecasting methods using the Nemenyi test on MASE values.
@@
Figure 6: Average ranks and 95% confidence intervals for all the forecasting methods using the Nemenyi test on CRPS values.
@@ 6.2 Managerial implications
Demand forecasting for contraceptives in developing countries is a critical managerial task, given the volatile and unpredictable nature of demand. However, many field-level staff still rely on basic methods like moving averages or demographic projections, which often fall short in addressing these complexities. Our research underscores the need to transition to advanced probabilistic forecasting approaches that provide a range of potential outcomes rather than a single-point estimate. This shift can enable field-level staff to better anticipate demand variability and uncertainty. For example, by using prediction intervals, field-level staff can align their decisions with specific priorities, such as minimizing stockouts by adopting more effective ordering strategies during periods of high variability.
@@ 6.3 Limitations and future directions
On the other hand, FPSC is often subject to uncertainties arising from complex demand patterns, variable lead times, and dependence on donor support [41]. For instance, demographic factors like the age structure of a region can influence contraceptive demand. Haakenstad et al. [26] highlighted that young women (ages 15-25) tend to prefer short-term contraceptive methods, while older, married women are more likely to use long-term methods. However, even these behaviors are heavily influenced by social and cultural beliefs [53]. In some regions, such as India, long-term contraceptive methods are popular among younger women [27]. Given such complexities, incorporating expert knowledge during the method-building process as predictive factors could help address these issues. Identifying and defining such influencing factors and incorporating them into forecasting methods remains an important research area for future exploration.
Another key issue in the FPSC is the presence of censored demand due to stockouts, under-reporting, or discontinuations. In our methoding process, we did not account for these scenarios. Future research should explore how to develop forecasting methods that can handle stockout data, mitigate its impact on decision-making. Moreover, addressing the challenges of cold starts (multiple origin points) and cold ends (discontinuations) in time series forecasting is crucial, as these are prevalent in FPSC and should be considered in future methods.

### v2

@@ 1 Introduction
To our knowledge, no previous work has focused on probabilistic forecasting in contraceptive demand estimation within the FPSC at the local healthcare site level. Thus, this paper first addresses this gap by investigating a probabilistic forecasting approach for estimating demand for contraceptive products using data from January 2016 to December 2019, extracted from the Logistics Management Information System (LMIS) of Cote d’Ivoire. This is the same dataset that was used in the competition. Moreover, since the publication of the Contraceptive Forecasting Handbook, which focuses on simple forecasting methods (see USAID [73] for more information), there has been no literature evaluating the applicability and usability of different forecasting methods in contraceptive demand forecasting. Therefore, in this study, we employ a range of forecasting methods, including time series, Bayesian, ML and foundational time series methods [^3] Foundational Time Series Methods: A class of pre-trained machine learning models designed specifically for time series forecasting, leveraging large-scale datasets to generalize across different forecasting tasks. These models require no or minimal tuning and can handle complex temporal patterns efficiently., to produce point forecasts along with probabilistic forecasts for all products across all healthcare sites.
@@
Given that system-generated forecasts and human forecasts offer distinct benefits, it is vital to design a hybrid intelligence system that combines them. In this context, where site-level demand planners produce point forecasts, we are particularly interested in how to combine point forecasts with probabilistic forecasts to produce combined probabilistic forecasts. However, the literature often treats point forecast combination methods and probabilistic forecast combination methods separately [77]. To address this gap, we propose a Constrained Quantile Regression Averaging (CQRA) method [^4] CQRA: A statistical method that combines multiple quantile regression models while enforcing constraints to ensure coherence and interpretability of predictions. It enhances robustness by optimizing weights assigned to different models under quantile-specific constraints, improving accuracy in heterogeneous data settings. to combine point forecasts made by experts with probabilistic forecasts generated by a system-based forecasting method. We compare the forecast performance using the Mean Absolute Scaled Error (MASE), a scale-independent metric that provides robustness and stability for point forecasts, and Continuous Ranked Probability Scores (CRPS), a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution for probabilistic forecasts in a cross-validation setup. Finally, we compare our method results against the submissions from the competition.
@@ 2.2 Human judgment in contraceptive demand forecasting
Various approaches exist for integrating human judgment with algorithmic forecasts, forming hybrid intelligence systems that leverage the strengths of both. Brau et al. [13] categorize these methods into five key types:
  • 1.
    Judgmental AdjustmentExperts modify algorithmic forecasts based on contextual insights.
  • 2.
    Quantitative CorrectionSystematic adjustments are applied to human forecasts using statistical techniques.
  • 3.
    Forecast CombinationSeparate judgmental and algorithm-based forecasts are merged into a single forecast.
  • 4.
    Judgment as a Model InputExpert knowledge is incorporated as a predictive variable within the forecasting model.
  • 5.
    Integrative Judgment LearningHuman inputs iteratively refine model predictions through a structured learning process.
Judgmental adjustments have been widely studied in supply chain contexts, particularly in promotional settings. Trapero et al. [71] found that structured human interventions during promotions improved forecasting accuracy. This aligns with findings from Fildes and Petropoulos [24], who argue that human adjustments enhance reliability when bias is minimized.
Quantitative correction methods offer a structured way to adjust forecasts while maintaining statistical integrity. Fildes et al. [23] highlight the importance of systematic corrections to mitigate bias and improve forecast accuracy.
Forecast combination has been explored as a way to merge expert forecasts with algorithmic outputs. Goodwin [28] highlights its effectiveness in capturing both statistical trends and domain knowledge, making it a useful technique when expert input is available alongside data-driven predictions.
Another structured approach to integrating expert knowledge into forecasting models is treating human judgment as a direct input to model-building. Instead of adjusting forecasts post hoc, this method incorporates expert-driven insights as predictive variables, allowing the model to factor in domain knowledge systematically. Arvan et al. [6] suggest that using expert judgment as a model input can improve forecasting accuracy, particularly in cases where statistical models alone struggle to capture contextual nuances. By embedding domain expertise within the forecasting process, this approach ensures that human insights are systematically integrated, rather than relying on manual interventions.
On the other hand, integrative Judgment Learning Baecke et al. [7], Brau et al. [13] treats human adjustments as predictive variables, allowing models to systematically weigh their impact. Similarly, Goodwin and Fildes [29] suggest that hybrid forecasting models should be designed to correct systematic biases in human judgment while maintaining adaptability.
Some studies explore hybrid forecasting methods that do not fit neatly into a single category. These approaches often involve adaptive weighting mechanisms optimization-driven model adjustments, or learning-based frameworks that dynamically integrate human and machine-generated forecasts [54]. While not explicitly categorized under the five methods, these strategies contribute to the broader understanding of how expert judgment and algorithmic forecasts can be effectively combined.
@@ 2.3 Literature limitations summary
Despite advancements in contraceptive demand forecasting, several key gaps remain unaddressed. First, the use of probabilistic forecasting in FPSC remains limited. While probabilistic methods have been widely explored in other domains, such as economic forecasting [26, 42], their application in FPSC is largely absent. Most existing studies rely on point forecasts, which do not account for demand uncertainty, making supply chain planning more vulnerable to unexpected fluctuations. Given the inherent unpredictability of contraceptive demand, incorporating probabilistic forecasting is critical for improving stock management and mitigating stockout risks
Second, contraceptive demand forecasting has primarily been conducted at the national or regional level [2, 37], often overlooking demand variability at individual healthcare sites. However, demand patterns vary significantly across locations [38], and without localized forecasts, procurement decisions may fail to align with actual site-level needs. This highlights the need for forecasting approaches that can adapt to site-specific demand while maintaining consistency with broader supply chain strategies.
Third, while forecast combination techniques have been extensively studied [58], their application in FPSC is underexplored. Existing studies typically assess individual forecasting methods without considering how integrating multiple forecasts could improve accuracy, particularly in healthcare supply chains. Additionally, there is limited empirical research comparing different forecasting methods for contraceptive demand estimation, making it difficult to determine the most effective approach in this context.
Finally, although judgmental adjustments are widely used in FPSC [23, 71], there is no structured approach that systematically integrates expert knowledge while ensuring coherence between human forecasts and probabilistic distributions in FPSC settings. In economic forecasting, Entropic Tilting has been used to refine probabilistic distributions based on external point forecasts [42, 46], yet these methods do not explicitly ensure coherence between expert judgment and forecast distributions. Our study addresses this gap by proposing a hybrid forecasting approach that systematically integrates expert forecasts with probabilistic models, ensuring that adjustments remain statistically grounded while incorporating expert insights.
@@ 3 Proposed hybrid approach
We propose a CQRA model to generate a combined probabilistic forecast, utilizing both point and probabilistic forecasts. This approach builds upon the CQRA model introduced by Wang et al. [78], which focuses on combining multiple probabilistic forecasts to produce a consolidated forecast distribution. The key concept in our proposed method is to generate quantiles from a given probabilistic forecast and adjust each quantile using weights. These weights are determined by treating the point forecast as the “new reality” and formulating a linear programming (LP) problem that minimizes both the pinball loss [^8] Pinball Loss: A proper scoring rule used to evaluate quantile forecasts, penalizing deviations based on whether the predicted quantile overestimates or underestimates the observed value. It ensures a well-calibrated probabilistic forecast by emphasizing accuracy across different quantile levels. and the absolute error between point forecast and mean of the weighted quantile forecast. The pinball loss is a strictly proper scoring rule used to evaluate quantile forecasts. It measures overall quantile performance by rewarding sharpness and penalizing miscalibration [34].
On the other hand, our proposed CQRA model shares similarities with the quantile combination approach of Trapero et al. [70] in that both methods determine weights by minimizing the tick-loss function. However, a fundamental difference is that while Trapero et al. [70] combine quantile forecasts generated from Kernel Density Estimators (KDE) and GARCH models to optimize safety stock levels, CQRA integrates expert point forecasts with probabilistic forecasts. This allows CQRA to adjust forecast distributions in response to human insights, particularly in settings where domain knowledge complements algorithmic predictions. By explicitly formulating an optimization problem that aligns expert-driven point forecasts with probabilistic distributions, CQRA provides a structured way to incorporate human intuition into data-driven forecasting models, making it well-suited for FPSC and other humanitarian supply chains.
On the other hand, Entropic Tilting has been used as a flexible approach to adjust predictive densities by incorporating external information while maintaining the statistical properties of the original model. This method systematically reweights the probability distribution of forecasts to ensure coherence with expert knowledge or external nowcasts. For example, Krüger et al. [42] applied Entropic Tilting to Bayesian VAR forecasts by incorporating short-term nowcasts, while Metaxoglou and Smith [46] used it to refine option-implied predictive densities for equity returns. These approaches allow for seamless adjustments of predictive distributions while minimizing distortions to the baseline statistical model. While Entropic Tilting provides a suitable method for combining probabilistic and point forecasts, our proposed CQRA model differs in its optimisation framework, which ensures alignment between expert-driven point forecasts and probabilistic forecasts in a constrained manner. Unlike Entropic Tilting, which reweights distributions without necessarily enforcing coherence between expert forecasts and the final probabilistic forecast, CQRA imposes constraints that ensure consistency between the adjusted forecast distribution and expert input. This makes CQRA particularly well-suited for supply chain applications where domain knowledge plays a critical role in refining model-based predictions.
@@
However, probabilistic forecasts may have a more dominant influence in this approach since the weights $w_{i}$ are restricted to the range $[0,1]$. This implies that point forecasts are expected to align closely with the forecast distribution. When the probabilistic forecasts are reliable and the point predictions do not significantly deviate from the mean of the probabilistic forecast, this approach is recommended. However, recognizing that this is often not the case in practice, we relax the sum-to-one constraint on the combination weights to provide greater flexibility in incorporating expert point forecasts and accounting for bias. A similar relaxation has been explored in prior research (e.g., [30]), which suggests that imposing weight constraints can sometimes lead to suboptimal forecasts by failing to account for systematic biases. Furthermore, Trapero et al. [70] also show that an unconstrained approach to optimizing quantile forecast combinations can yield improvements, particularly in applications requiring adaptive weight adjustments. By allowing weights to exceed unity and introducing a bias factor, our modified CQRA approach accommodates situations where expert forecasts systematically deviate from model-based predictions, enhancing the adaptability and robustness of the combined forecast.
@@ 4.2 Data exploration
Figure 4: Trend strength and seasonality in the time series of stock distribution. Each point in the scatter plot represents one of the 1,360 time series analysed, with trend and seasonality strengths measured on a scale from 0 to 1 (0 indicating weak and 1 indicating strong).
@@ 4.3 Forecasting setup
In our cross-validation setup, we define the training period as all available data up to September 2019, ensuring sufficient historical observations for model learning. The testing period consists of rolling evaluation windows, where each test set spans a 3-step-ahead forecasting horizon to align with the competition requirements. At each iteration, the training set expands while the test set moves forward by one step, maintaining a consistent evaluation structure. We limited the number of rolling origins to 3 per series due to computational constraints, but this still provided us with meaningful insights into method performance over time. For forecasting, we employed recursive multi-step forecasting and we generated 1000 future paths per a series. All method development and hyper-parameter tuning were conducted using only the training data to ensure that the evaluation remained unbiased and the methods were properly validated.
@@ 4.3.1 Probabilistic forecasting using bootstraping
Given that our study employs multiple forecasting methods, bootstrapping provides a unified framework that can be applied consistently without requiring additional parametric assumptions. Moreover, it allows us to approximate the empirical distribution of forecast errors without imposing strict distributional constraints, making it particularly suitable for datasets with varying demand patterns in the FPSC.
@@ 4.5 Performance evaluation
To assess the performance of our forecasting methods, we used both point forecast and probabilistic forecast evaluation metrics. We evaluated point forecasts using the MASE. MASE was chosen for two primary reasons. First, it was the official evaluation metric used in the USAID Intelligent Forecasting Competition, allowing us to directly compare our model performance with previous benchmark results. Second, MASE is a scale-independent metric that provides robustness, and stability [41].
The MASE formula is:
@@
To evaluate the accuracy of probabilistic forecasts, we employed the CRPS, a widely used metric in probabilistic forecasting that assesses the sharpness and calibration of the forecast distribution.
The CRPS is given by:
@@
While CRPS provides a comprehensive evaluation of the entire predictive distribution, there are cases where accuracy at specific quantiles is of particular interest. For example, in inventory management of family planning health commodities, higher quantiles (e.g., the 95th percentile) might be important to ensure efficient stock management and maintain a high service level. By accurately capturing demand at these upper quantiles, supply chain planners can better mitigate stockouts and ensure the consistent availability of essential health supplies. In such scenarios, the quantile score (or pinball score) becomes a more appropriate metric, as it directly evaluates forecast accuracy at the chosen quantile, ensuring a more targeted assessment of predictive performance.
@@ 5 Analysis and results
Figure 5: Average ranks of forecasting methods with 95% confidence intervals based on the Nemenyi test for MASE values. Lower ranks indicate better forecast performance.
@@
Figure 6: Average ranks of forecasting methods with 95% confidence intervals based on the Nemenyi test for CRPS values. Lower ranks indicate better forecast performance.
@@ 6.2 Managerial implications
Demand forecasting for contraceptives in developing countries is a critical managerial task, given the volatile and unpredictable nature of demand. However, many field-level staff still rely on basic methods like moving averages or demographic projections, which often fall short in addressing these complexities. Our research underscores the need to transition to advanced probabilistic forecasting approaches that provide a range of potential outcomes rather than a single-point estimate. This shift can enable field-level staff to better anticipate demand variability and uncertainty. Thus, our proposed model can support procurement strategies by providing probabilistic forecasts that help optimize order quantities based on demand uncertainty. This enables field level staff to reduce both overstocking and understocking risks by quantifying the uncertainty. Additionally, the model enhances FPSC resilience by allowing inventory levels to be adjusted dynamically based on forecast distributions, improving responsiveness to demand fluctuations while ensuring consistent contraceptive availability.
@@ 6.3 Limitations and future directions
On the other hand, FPSC is often subject to uncertainties arising from complex demand patterns, variable lead times, and dependence on donor support [49]. For instance, demographic factors like the age structure of a region can influence contraceptive demand. Haakenstad et al. [31] highlighted that young women (ages 15-25) tend to prefer short-term contraceptive methods, while older, married women are more likely to use long-term methods. However, even these behaviours are heavily influenced by social and cultural beliefs [62]. In some regions, such as India, long-term contraceptive methods are popular among younger women [32]. Infrastructure variability also poses challenges, particularly in rural and underserved areas, where logistical constraints can affect the distribution and accessibility of contraceptive products. Furthermore, Karimi et al. [38] highlight that rural facilities often face difficulties such as poor road conditions, inadequate storage, and delivery delays, further complicating supply chain operations. While field-level healthcare staff rely on their contextual knowledge to adjust forecasts, much of this information remains undocumented, making it difficult to systematically incorporate into forecasting models. Identifying and defining such influencing factors and integrating them into forecasting methods remains an important research area for future exploration.
Another key issue in the FPSC is the presence of censored demand due to stockouts, under-reporting, or discontinuations. In our modelling process, we did not account for these scenarios. Future research should explore how to develop forecasting methods that can handle stockout data, mitigate its impact on decision-making. Moreover, addressing the challenges of cold starts (multiple origin points) and cold ends (discontinuations) in time series forecasting is crucial, as these are prevalent in FPSC and should be considered in future methods.
@@
Another limitation of our study is we did not conduct an exploratory analysis of the sources of expert bias in judgemental forecasting. However, understanding when and why human forecasters introduce biases is crucial for improving hybrid forecasting accuracy. Future research could address this gap by systematically analysing the conditions under which expert bias occurs, particularly in the FPSC context. This could involve access to detailed expert forecasts, identifying systematic biases, and developing mechanisms to mitigate their impact. Such insights would further refine hybrid intelligence models by improving the integration of human intuition and algorithmic precision.